# Towards Interoperable Data Spaces: Comparative Analysis of Data Space Implementations between Japan and Europe

### v1

@@ Abstract
The rapid evolution of data spaces is transforming the landscape of secure and interoperable data sharing across industries and geographies. In Europe, the concept of data spaces, supported by initiatives such as the European Data Strategy, emphasises the importance of trust, sovereignty, and interoperability. Meanwhile, Japan has been developing its approach to data sharing, in line with global trends but also to address unique domestic challenges. Despite these parallel advances, achieving interoperability between European and Japanese data spaces remains a critical challenge due to differences in governance, technology standards, and authentication frameworks. This paper undertakes a comparative analysis of DATA-EX and Catena-X to explore the challenges and opportunities for achieving interoperability between Japanese and European data spaces. By examining common data exchange processes, key objects such as participants, datasets, and data catalogs, and specific evaluation criteria, the study identifies gaps and proposes actionable solutions. Through this analysis, the paper aims to contribute to the ongoing discourse on global data interoperability. It proposes an interoperable architecture that bridges regional differences while addressing common challenges. It also identifies challenges that should be addressed to achieve interoperability.
@@ 1 INTRODUCTION
The rapid evolution of data spaces is transforming the landscape of secure and interoperable data sharing across industries and regions. In Europe, the concept of data spaces, supported by initiatives such as the European Data Strategy, emphasises the importance of trust, sovereignty, and interoperability. Meanwhile, Japan has been developing its approach to data sharing, in line with global trends but also addressing unique domestic challenges. Despite these parallel advancements, achieving interoperability between European and Japanese data spaces remains a critical challenge due to differences in governance, technology standards, and authentication frameworks.
This paper undertakes a comparative analysis of DATA-EX and Catena-X to explore the challenges and opportunities for achieving interoperability between Japanese and European data spaces. By examining common data exchange processes, key objects such as participants, datasets, and data catalogs, and specific criteria for evaluation, the study identifies gaps and proposes actionable solutions. The findings are based on a thorough analysis of relevant documentation, highlighting areas for alignment and improvement.
Through this analysis, the paper aims to contribute to the ongoing discourse on global data interoperability. It proposes an interoperable architecture that bridges regional differences while addressing common challenges.
The rest of the paper is structured as follows. Section 2 reviews the development of data spaces in Europe and Japan, and examines the role of trust anchors and interoperability challenges. Section 3 describes the analysis method, including key evaluation criteria and data sources. Section 4 presents the analysis results, comparing participants, devices, datasets, and other key elements of the two platforms. Section 5 discusses identified gaps, proposes an interoperable architecture, and addresses specific challenges. Section 6 reviews related works, and Section 7 concludes with findings and future directions.
@@ 2.1 Development of Data Spaces in Europe
Europe has been accelerating its efforts to maximise the potential of the data economy. Central to these efforts are the European Data Strategy [^1] https://commission.europa.eu/strategy-and-policy/priorities-2019-2024/europe-fit-digital-age/european-data-strategy˙en and the concept of data sovereignty. The European Data Strategy aims to foster data sharing within and beyond the region, positioning Europe as a global leader in the data economy. This strategy is rooted in the belief that data should be managed and utilised fairly, without being disproportionately controlled by a few companies or nations. Data sovereignty refers to the principle that entities owning and managing data have the exclusive right to determine its usage. For example, this includes the control over the data, even if the data is shared with other parties. Europe has reinforced this principle through stringent data protection regulations such as General Data Protection Regulation (GDPR), ensuring that individuals and organisations can safeguard their rights while contributing to the data economy. Initiatives like International Data Spaces Association (IDSA) and Gaia-X exemplify Europe’s approach to data spaces. IDSA focuses on creating secure and standardised frameworks for data sharing across various industries, ensuring that data sovereignty is maintained. Meanwhile, Gaia-X, a European initiative for cloud infrastructure and data sharing, aims to establish an interoperable ecosystem where data and services can be securely exchanged while adhering to European values such as transparency and privacy. Together, these initiatives highlight Europe’s commitment to fostering a trusted data economy by addressing technical, legal, and ethical challenges inherent in cross-border data exchanges. With this multi- pronged strategy, Europe seeks to establish itself as a leader in the development of a robust, scalable, and ethical data ecosystem.
@@ 2.2 Development of Data Spaces in Japan
Similarly, Japan has pursued its own approach to developing data spaces, centerd on the concepts of Society 5.0222https://www8.cao.go.jp/cstp/english/society5˙0/ and Data Free Flow with Trust (DFFT) [^3] https://www.digital.go.jp/en/policies/dfft. Society 5.0 is a national strategy that envisions a super-smart society where cyberspace and physical space are seamlessly integrated to address societal challenges through data utilisation. Within this concept, DFFT emphasises the importance of trust in data flow, enabling secure data sharing across national and organisational boundaries. The second phase of the Strategic Innovation Promotion Program (SIP) and its subsequent evolution into DATA-EX [Koshizuka and Mano, 2022] have constituted significant advances in Japan’s data space endeavours, spearheaded by governmental initiatives. In the context of SIP, a connector called Connector Architecture for Decentralized Data Exchange (CADDE) [^4] https://sip-cyber-x.jp/en/ has been developed with the objective of enabling secure and efficient data sharing between organisations. This foundational technology constituted the basis for the subsequent establishment of the DATA-EX initiative, which was launched within Data Society Alliance (DSA). The objective of this initiative is to expand the capabilities of CADDE into a broader data exchange infrastructure, thereby promoting interoperability and trust in data sharing across sectors. Another notable effort is the Ouranos Ecosystem[Ministry of Economy, Trade and Industry, 2023] project, which began in 2022 under the leadership of Information Processing Agency’s Digital Architecture Design Center (IPA DADC). The project’s objective is to enhance battery traceability within the automotive manufacturing sup- ply chain as well as ensure compliance with European battery regulations. This addresses the critical need for transparency and accountability in tracking battery production and usage. In 2023, the implementation phase was completed, establishing a technical foundation for real-world application. By 2024, the operational body, Automotive and Battery Traceability Center Association (ABtC), was established, marking the transition to full-scale operation and contributing to the advancement of sustainable supply chain management.
@@ 2.4 Challenges of Interoperability Between Japan and Europe
Currently, data spaces in Japan and Europe are built on different technological bases, making interoperability challenging. In Europe, standardisation of data space and trust is being pursued in IDSA and Gaia-X respectively. In Japan, meanwhile, the DATA-EX and Ouranos Ecosystem are pursuing their own individual standardisation agendas. However, the need for interoperability between these data spaces is becoming increasingly critical, particularly for addressing global issues such as carbon footprint transparency and international supply chain management. A significant challenge lies in ensuring the trustworthiness of participating entities and the data exchanged. Without robust mechanisms to guarantee trust, resistance to data sharing may increase, stalling the flow of data and hindering the data economy. Therefore, creating systems that integrate or harmonise the technological foundations of Japan and Europe is essential. Addressing this challenge will pave the way for a sustainable and interconnected data economy in the future.
@@ 3 ANALYSIS METHOD
As a first step towards solving the challenge mentioned above, we undertake a comparative analysis of two prominent data space platforms in Japan and Europe from a trust perspective. We selected DATA-EX in Japan and Catena-X in Europe as the representative data space platforms due to substantial number of analysable documents and past maturity evaluation which is conducted in [Dam et al., 2023] and [Bacco et al., 2024].
@@ 3.2 Target Objects
Device (server and endpoint which connector works)
@@ 3.4 Documents Used for Analysis
With regard to Catena-X, we selected technical documents in both the technical standards [Catena-X, 2024] and deliverables of the Eclipse Tractus-X project [^7] https://github.com/eclipse-tractusx. Catena-X Automotive Network e.V. is publishing technical standards, but it doesn’t cover all specifications of Catena-X. Eclipse Tractus-X is an open-source project to develop reference implementations of Catena-X and providing some technical documents (not comprehensive) and source code. The version of the selected documents was adhered to “CX-Jupiter” release as Catena-X is under continuous development.
@@ 5 DISCUSSION
Based on the analysis results above, we discuss the challenges of the international connectivity of data spaces. First, we review the gaps identified as shown in Table 4. Second, we discuss an interoperable architecture based on the review. Finally, we discuss two challenges based on the interoperable architecture.
@@ 5.1.6 Sending & Receiving Log
5.2 Interoperable Architecture
Figure 5 shows the possible interoperable architecture between DATA-EX and Catena-X based on our gap analysis.
Figure 5: Interoperable architecture between DATA-EX and Catena-X.
To discuss interoperable architecture, we judged each gap whether it interferes interoperability or not. For an object which has the former gap, we thought that we should use the function or method on each platform as they are. In our judgement, only the participant has the former gap, so a connector, which is a component to verify the participant, and an account, which acts as a participant’s credential, should be prepared in both platforms. Fast and economic way to prepare the connector and account will be desired because there is large burden in preparation against two platforms.
@@ 5.4 A Challenge on The Dataset
To address this problem, we consider two solutions. The first solution is to create an international repository. Although it is an ideal solution, it takes a long time to proceed mutual recognition between countries or regions. The second solution is to create an index repository to show correspondence of semantics among existing repository in each data space. It is more realistic than the first solution because registration and maintenance of the index can be led by the participant. Automation of registration or maintenance using technologies such as a name identification method is also suitable for this solution. Thus, we believe the second solution should be proceeded ahead of the first one.
@@ 6 RELATED WORKS
A basic survey against data space connectors was carried out by [Dam et al., 2023] and [Bacco et al., 2024]. A total of eight connectors including EDC used in Catena-X were surveyed and evaluated, but Japanese connectors are not included in the surveyed connectors. Moreover, it didn’t discuss the interoperability issues between connectors, while [Bacco et al., 2024] explains IDS RAM[IDSA, 2023] as a common specification for data spaces. Our work is to compare European connectors in accordance with IDS RAM and Japanese connectors not in accordance with IDS RAM and to consider interoperability among them.
@@
Attempts to create design options for data spaces by analysis of published papers or interview were conducted by [Gieß et al., 2023] and [Gieß et al., 2025]. This paper also conducted a frequency analysis of design options against existing data spaces. Their design options include a perspective of trust. They emphasise the dataset and log as objects to consider trust, however, many other objects, such as contracts, are generated in an actual data exchange process. Our work is to conduct an analysis with comprehensive criteria covering a whole data exchange process.
@@
There are two papers that consider interoperability between data spaces. The first is [Schinke and Roßmann, 2024]. They investigated and analysed various technologies (including the data space) that can be used to increase data sharing. They proposed six ways to enable trustworthy data sharing using the technologies they researched. One of six ways included an architecture that combines multiple data spaces. They stated that handling of a dataset for multiple data spaces is a challenge in this architecture and this challenge should be solved by data trustee[Schinke et al., 2023] [Steinert et al., 2024]. The data trustee can hide a layer of connectors from the participants by creating intermediary layer, so the participant does not have to consider interoperability issues. While our work is to discuss an interoperable architecture without any trust intermediary like the data trustee, the data trustee can be another solution towards an interoperable architecture.
@@
The second one is [Hutterer and Krumay, 2024]. They surveyed and analysed the scopes of governance in data spaces from published papers. As a result of the analysis, three types of governance were identified: ecosystem governance, technological governance, and operational governance. In the technological governance, the authors argues that the federated data space architecture[Schleimer et al., 2023] can achieve interoperability across different data spaces. The federated data space architecture expresses the architecture to realise global data availability across multiple data spaces as three logical layers, local foundation layer, federation layer, and global presentation layer. This architecture was evaluated with a circular economy example, but designs for each layer were abstract because no actual connectors were described. Our work is to discuss an interoperable architecture at a more concreate level by comparing and analysing actual connectors.
@@
An experimental environment (called as International Testbed for Dataspace Technology, or ITDT) to try interoperability measures was proposed by [Matsunaga et al., 2023]. The ITDT aims to enable researchers or developers to test connectivity or interoperability among connectors which built on different technological bases. The ITDT can be an adequate environment for us to implement and test our proposals in future, although a DATA-EX connector has not provided yet in the ITDT.

### v2

@@ Abstract
Data spaces are evolving rapidly. In Europe, the concept of data spaces, which emphasises the importance of trust, sovereignty, and interoperability, is being implemented as a platform such as Catena-X. Meanwhile, Japan has been developing its approach to data sharing, in line with global trends but also to address unique domestic challenges, resulting a platform such as DATA-EX. Achieving interoperability between European and Japanese data spaces remains a critical challenge due to the differences created by these parallel advances. Although interoperability between data spaces has several aspects, compatibility of trust in the participating entities and the data exchanged is a significant aspect due to its influence on business. This paper undertakes a comparative analysis of DATA-EX and Catena-X while focusing on aspect of trust, to explore the challenges and opportunities for achieving interoperability between Japanese and European data spaces. By examining common data exchange processes, key objects such as datasets, and specific evaluation criteria, the study identifies gaps, challenges, and proposes actionable solutions such as inter-exchangeable topology. Through this analysis, the paper aims to contribute to the ongoing discourse on global data interoperability.
@@ 1 INTRODUCTION
The rapid evolution of data spaces is transforming the landscape of secure and interoperable data sharing across industries and regions. In Europe, the concept of data spaces, supported by initiatives such as the European Data Strategy, emphasises the importance of trust, sovereignty, and interoperability. Meanwhile, Japan has been developing its approach to data sharing, in line with global trends but also addressing unique domestic challenges. Despite these parallel advancements, achieving interoperability between European and Japanese data spaces remains a critical challenge due to differences in governance, technology standards, and authentication frameworks. Interoperability between data spaces has various aspects such as compatibility of technologies like connectors, compatibility of objects handled in data spaces, and compatibility of trust on participating entities and the data exchanged. Compatibility of trust is a significant issue among these aspects, because actual data exchange is never established without trustworthiness between participants.
This paper undertakes a comparative analysis of DATA-EX and Catena-X to explore the challenges and opportunities for achieving interoperability between Japanese and European data spaces. The analysis is focusing on trust. By examining common data exchange processes, key objects such as participants, datasets, and data catalogs, and specific criteria for evaluation, the study identifies gaps and proposes actionable solutions. The findings are based on a thorough analysis of relevant documentation, highlighting areas for alignment and improvement.
Through this analysis, the paper aims to contribute to the ongoing discourse on global data interoperability. It proposes an inter-exchangeable topology that bridges regional differences while addressing common challenges.
The rest of the paper is structured as follows. Section 2 reviews the development of data spaces in Europe and Japan, and examines the role of trust anchors and interoperability challenges. Section 3 describes the analysis method, including key evaluation criteria and data sources. Section 4 presents the analysis results, comparing participants, devices, datasets, and other key elements of the two platforms. Section 5 discusses identified gaps, proposes an inter-exchangeable topology, and addresses specific challenges. Section 6 reviews related works, and Section 7 concludes with findings and future directions.
@@ 2.1 Development of Data Spaces in Europe
Europe has been accelerating its efforts to maximise the potential of the data economy. Central to these efforts are the European Data Strategy [^1] https://commission.europa.eu/strategy-and-policy/priorities-2019-2024/europe-fit-digital-age/european-data-strategy˙en and the concept of data sovereignty. The European Data Strategy aims to foster data sharing within and beyond the region, positioning Europe as a global leader in the data economy. This strategy is rooted in the belief that data should be managed and utilised fairly, without being disproportionately controlled by a few companies or nations. Data sovereignty refers to the principle that entities owning and managing data have the exclusive right to determine its usage. For example, this includes the control over the data, even if the data is shared with other parties. Europe has reinforced this principle through stringent data protection regulations such as General Data Protection Regulation (GDPR), ensuring that individuals and organisations can safeguard their rights while contributing to the data economy. Initiatives like International Data Spaces Association (IDSA) and Gaia-X exemplify Europe’s approach to data spaces. IDSA focuses on creating secure and standardised frameworks for data sharing across various industries, ensuring that data sovereignty is maintained. Meanwhile, Gaia-X, a European initiative for cloud infrastructure and data sharing, aims to establish an interoperable ecosystem where data and services can be securely exchanged while adhering to European values such as transparency and privacy. Together, these initiatives highlight Europe’s commitment to fostering a trusted data economy by addressing technical, legal, and ethical challenges inherent in cross-border data exchanges. With this multipronged strategy, Europe seeks to establish itself as a leader in the development of a robust, scalable, and ethical data ecosystem.
@@ 2.2 Development of Data Spaces in Japan
Similarly, Japan has pursued its own approach to developing data spaces, centerd on the concepts of Society 5.0222https://www8.cao.go.jp/cstp/english/society5˙0/ and Data Free Flow with Trust (DFFT) [^3] https://www.digital.go.jp/en/policies/dfft. Society 5.0 is a national strategy that envisions a super-smart society where cyberspace and physical space are seamlessly integrated to address societal challenges through data utilisation. Within this concept, DFFT emphasises the importance of trust in data flow, enabling secure data sharing across national and organisational boundaries. The second phase of the Strategic Innovation Promotion Program (SIP) and its subsequent evolution into DATA-EX [Koshizuka and Mano, 2022] have constituted significant advances in Japan’s data space endeavours, spearheaded by governmental initiatives. In the context of SIP, a connector called Connector Architecture for Decentralized Data Exchange (CADDE) [^4] https://sip-cyber-x.jp/en/ has been developed with the objective of enabling secure and efficient data sharing between organisations. This foundational technology constituted the basis for the subsequent establishment of the DATA-EX initiative, which was launched within Data Society Alliance (DSA). The objective of this initiative is to expand the capabilities of CADDE into a broader data exchange infrastructure, thereby promoting interoperability and trust in data sharing across sectors. Another notable effort is the Ouranos Ecosystem [Ministry of Economy, Trade and Industry, 2023] project, which began in 2022 under the leadership of Information Processing Agency’s Digital Architecture Design Center (IPA DADC). The project’s objective is to enhance battery traceability within the automotive manufacturing supply chain as well as ensure compliance with European battery regulations. This addresses the critical need for transparency and accountability in tracking battery production and usage. In 2023, the implementation phase was completed, establishing a technical foundation for real-world application. By 2024, the operational body, Automotive and Battery Traceability Center Association (ABtC), was established, marking the transition to full-scale operation and contributing to the advancement of sustainable supply chain management.
@@ 2.4 Challenges of Interoperability Between Japan and Europe
Currently, data spaces in Japan and Europe are built on different technological bases, making interoperability challenging. In Europe, standardisation of data space and trust is being pursued in IDSA and Gaia-X respectively. In Japan, meanwhile, the DATA-EX and Ouranos Ecosystem are pursuing their own individual standardisation agendas. However, the need for interoperability between these data spaces is becoming increasingly critical, particularly for addressing global issues such as carbon footprint transparency and international supply chain management. Interoperability between data spaces has various aspects such as compatibility of technologies which means connectors or related components can connect each other seamlessly beyond a data space, compatibility of objects which means expression of objects handled in data spaces can be interpreted uniquely, and compatibility of trust which means trustworthiness of participating entities and the data exchanged is similar among data spaces. A significant challenge lies in the compatibility of trust. Without robust mechanisms to guarantee trust, resistance to data sharing may increase, stalling the flow of data and hindering the data economy. Therefore, creating systems that integrate or harmonise the technological foundations of Japan and Europe is essential. Addressing this challenge will pave the way for a sustainable and interconnected data economy in the future.
@@ 3 ANALYSIS METHOD
As a first step towards solving the challenge mentioned above, we undertake a comparative analysis of two prominent data space platforms in Japan and Europe from a trust perspective. We selected DATA-EX in Japan and Catena-X in Europe as the representative data space platforms due to substantial number of analysable documents and the past maturity evaluation [Dam et al., 2023] [Bacco et al., 2024].
@@ 3.2 Target Objects
Device (server and endpoint on which connector works)
@@ 3.4 Documents Used for Analysis
With regard to Catena-X, we selected technical documents in both the technical standards [Catena-X, 2024] and deliverables of the Eclipse Tractus-X project [^7] https://github.com/eclipse-tractusx. Catena-X Automotive Network e.V. is publishing technical standards, but it doesn’t cover all specifications of Catena-X. Eclipse Tractus-X is an open-source project to develop reference implementations of Catena-X and providing some technical documents (not comprehensive) and source code. The version of the selected documents was adhered to CX-Jupiter release as Catena-X is under continuous development.
@@ 5 DISCUSSION
Based on the analysis results above, we discuss the challenges of the international connectivity of data spaces. First, we review the gaps identified as shown in Table 4. Second, we discuss an inter-exchangeable topology based on the review. Third, we discuss two challenges based on the topology. Finally, we discuss possible policy alignment for addressing two challenges.
@@ 5.1.6 Sending & Receiving Log
5.2 Inter-exchangeable Topology
Figure 5 shows the possible inter-exchangeable topology between DATA-EX and Catena-X based on our gap analysis.
Figure 5: Inter-exchangeable topology between DATA-EX and Catena-X.
To discuss inter-exchangeable topology, we judged each gap whether it interferes interoperability or not. For an object which has the former gap, we thought that we should use the function or method on each platform as they are. In our judgement, only the participant has the former gap, so a connector, which is a component to verify the participant, and an account, which acts as a participant’s credential, should be prepared for both platforms like a dual-stack interface. Fast and economic way to prepare the connector and account will be desired because there is large burden in preparation against two platforms.
@@ 5.4 A Challenge on The Dataset
To address this problem, we consider two solutions. The first solution is to create an international repository. Although it is an ideal solution, it takes a long time to proceed mutual recognition between countries or regions. The second solution is to create an index repository to show correspondence of semantics among existing repository in each data space. It is more realistic than the first solution because registration and maintenance of the index can be led by the participant. To implement index repository, some existing semantic technologies such as knowledge graphs [Theissen-Lipp et al., 2023] may be useful. Automation of registration or maintenance using technologies such as a name identification method is also suitable for this solution. Thus, we believe the second solution should be proceeded ahead of the first one.
5.5 Possible Policy Alignment
For address two challenges above in an ideal way, policy alignment and mutual recognition between countries or regions is essential. There are two approaches to achieving it. The first is standardization. While the IDSA standardise the data space itself and related technologies in the data space protocol, Gaia-X standardise trust required to join the data space ecosystem in the Gaia-X compliance document [Gaia-X, 2024]. The Gaia-X compliance document defines trust standards to be satisfied as some levels and laws or regulations to be followed to be certified at each level. Gaia-X itself is a European organization, so approach to develop trust standards applicable internationally through international and neutral standardization organizations is necessary.
The second approach is discussion in the international committee. At G7 Hiroshima 2023, the establishment of the Institutional Arrangement for Partnership (IAP)  [^9] https://www.digital.go.jp/en/policies/dfft/dfft-iap was endorsed by G7 leaders to operationalise DFFT. The IAP can provide opportunities for discussion utsing existing committees or organs of international organisations, so discussing challenges here can be another approach.
@@ 6 RELATED WORKS
A basic survey against data space connectors was carried out [Dam et al., 2023] [Bacco et al., 2024]. A total of eight connectors including EDC used in Catena-X were surveyed and evaluated, but Japanese connectors are not included in the surveyed connectors. Moreover, it didn’t discuss the interoperability issues between connectors, while they explains IDS RAM [IDSA, 2023] as a common specification for data spaces. Our work is to compare European connectors in accordance with IDS RAM and Japanese connectors not in accordance with IDS RAM and to consider interoperability among them.
@@
Attempts to create design options for data spaces by analysis of published papers or interview were conducted [Gieß et al., 2023] [Gieß et al., 2025]. This paper also conducted a frequency analysis of design options against existing data spaces. Their design options include a perspective of trust. They emphasise the dataset and log as objects to consider trust, however, many other objects, such as contracts, are generated in an actual data exchange process. Our work is to conduct an analysis with comprehensive criteria covering a whole data exchange process.
@@
Consideration of interoperability between data spaces was carried out [Schinke and Roßmann, 2024]. They investigated and analysed various technologies (including the data space) that can be used to increase data sharing. They proposed six ways to enable trustworthy data sharing using the technologies they researched. One of six ways included an architecture that combines multiple data spaces. They stated that handling of a dataset for multiple data spaces is a challenge in this architecture and this challenge should be solved by data trustee [Schinke et al., 2023] [Steinert et al., 2024]. The data trustee can hide a layer of connectors from the participants by creating intermediary layer, so the participant does not have to consider interoperability issues. While our work is to discuss an inter-exchangeable topology without any trust intermediary like the data trustee, the data trustee can be another solution towards achievement of interoperability.
@@
Another consideration of interoperability between data spaces was carried out [Hutterer and Krumay, 2024]. They surveyed and analysed the scopes of governance in data spaces from published papers. As a result of the analysis, three types of governance were identified: ecosystem governance, technological governance, and operational governance. In the technological governance, the authors argues that the federated data space architecture [Schleimer et al., 2023] can achieve interoperability across different data spaces. The federated data space architecture expresses the architecture to realise global data availability across multiple data spaces as three logical layers, local foundation layer, federation layer, and global presentation layer. This architecture was evaluated with a circular economy example, but designs for each layer were abstract because no actual connectors were described. Our work is to discuss an inter-exchangeable topology at a more concreate level by comparing and analysing actual connectors.
@@
An experimental environment (called as International Testbed for Dataspace Technology, or ITDT) to try interoperability measures was proposed [Matsunaga et al., 2023]. The ITDT aims to enable researchers or developers to test connectivity or interoperability among connectors which built on different technological bases. The ITDT can be an adequate environment for us to implement and test our proposals in future, although a DATA-EX connector has not provided yet in the ITDT.